# HF radar estimation of ocean wave parameters: second-order Doppler spectrum versus Bragg wave modulation approach

### v1

@@ 4.3 Omnidirectional estimation
Figure 13: Scatter plot of omnidirectional radar $H_{s}$ from (a and c) SODS method and (b and d) FOBP method versus $H_{s}$ from CANDHIS buoy where each point is colored by the spatial density of nearby points. The spatial locations are along the ranges $8$ (a and b) first row, and $20$ (c and d) second row.
@@ 4.4 Comparison with the literature
A number of studies have been published in the literature on the estimation of wave heights and periods from a single HFR station and a single radar frequency. Some of them reported systematic comparisons with in situ instruments over a significantly long period (at least several months) using either phased array (Wyatt et al.,, 2003; Ramos et al.,, 2009; Gomez et al.,, 2015; Lopez et al.,, 2016; Lopez and Conley,, 2019; Alattabi et al.,, 2019; Al-Attabi et al.,, 2021) or compact radar systems (Roarty et al.,, 2015; Zhou and Wen,, 2015; Long et al.,, 2011; Atan et al.,, 2016; Saviano et al.,, 2020; Basañez et al.,, 2020). All of these methods are based on the exploitation of the second-order Doppler spectrum using either Barrick’s original methods, the full wave spectrum inversion method, or empirical methods developed by the radar manufacturers. The performance obtained varies considerably depending on the geographical location, the radar frequency, the measurement range, the antenna array, the inversion method chosen, the data post-processing choices, etc. However, two general trends can be identified within this diversity of results. First, the estimation of $H_{s}$ with phased array systems is generally slightly better than that with compact systems, with high Pierson correlation coefficients ($>0.9$) and RMSE often in the order of 30-40 cm and less than about 60 cm in all studies. The observed RMSE with compact systems are generally larger, in the order of 40-80 cm, with a large scatter of values and in some cases very large RMSE ($>1$ m); the associated correlation coefficients are generally smaller, mostly in the range of 0.5-0.9, except for the study Atan et al., (2016), which reports exceptionally good performances of 18-29 cm RMSE and 0.78-0.86 Pierson correlation coefficient. However, the decision to eliminate the $>25$% outliers (when compared with buoy measurement) before calculating the statistics is questionable. The second point is that the estimation of $T_{p}$ is less systematic and generally of lower quality than that of $H_{s}$, with RMSE of the order of $1-3$ seconds (to be compared with the usual range of periods, say $6-10$ seconds) and Pierson correlation coefficients of the order of 0.5-0.7.
Our results are in agreement with these trends. The RMSE obtained for the estimation of $H_{s}$ with SODS ranges from $0.39$ to about $0.5$ m for a wide angular sector in the near and medium range (up to about 20 km, see Figure 8c) and is slightly larger in the omnidirectional case, which is more representative of a compact antenna array. The estimation of $T_{p}$ is much coarser, with a lower correlation coefficient and an RMSE of at best $1.3$ seconds (that is a relative error of the order of $25$%). Note that we could have improved the statistical comparisons by additional post-processing of the data (elimination of outliers, jumps, etc.), but we chose to limit the impact of post-processing, which can be misleading in the evaluation of the method.
@@ 5 Conclusion
In this work, we have tested the accuracy of the classical Barrick’s method for estimating the main sea state parameters based on weighted integrals of the second-order backscattered Doppler spectrum. For this purpose, we analyzed a one-year HFR data set from a monostatic station in the Toulon region. In this region, the wave characteristics are homogeneous, allowing us to use both the directional and omnidirectional signals to estimate the significant wave height and peak period within the radar coverage area. The results presented in this study have been compared with the data collected by the CANDHIS buoy located near the island of Porquerolles. These data from the buoy have been considered as the ground truth for the analyses presented in this study. For the classical Barrick’s method with directional radar signal, the results obtained for significant wave height have a good accuracy compared to the time series of the CANDHIS buoy, with a very high correlation and a low RMSE value between both of them (close to $90\%$ and around $0.4$m, respectively, up to $30$km from the radar). Regarding the estimation of the wave peak period, this method shows some deficiencies in capturing low period values. This may be due to the limitation of differentiating the second-order spectrum from the signal noise or to a lower precision in the asymptotic formula of the weighting function at low wind speeds. The main limitations of this method are its sensitivity to electromagnetic interference and its degradation as we move further from the radar, reaching a point, in this study about $30$ km from the radar, where the results are unreliable.
In addition, a new method based on the modulation of the first order of the backscattered Doppler spectrum was presented in this study. This method allows wave characteristics to be estimated at greater distances from the radar since the first order of the spectrum is generally much more energetic than the second order and is easier to distinguish at great distances. The results obtained for the same year selected for Barrick’s method using the directional signal of the radar show that this new method gives a good quality in the estimation of the significant wave height, obtaining a better coverage in the estimation (up to 52.5 km compared to 30 km for the Barrick’s method). As for the estimation of the peak period, this new method gives better results than Barrick’s method even at distances close to the radar, with correlations close to $80\%$ and an RMSE of $1.5$s. While this method gives good results in estimating the peak period of the wave, it underestimates the buoy values for small periods, which requires some improvement in future work. The main limitation of this new method so far is that, at least for the $H_{s}$, it requires a preliminary calibration procedure with some ground truth measurement. This is not the case for the SODS method, which is based on the ratio of the first and second-order Doppler spectra and is therefore insensitive to absolute calibration.
In this study, in addition to treating the directional radar signal to estimate the wave characteristics, omnidirectional signals have also been evaluated to test the quality of both methods when it is impossible to perform the beamforming technique and to distinguish from which azimuth the signal arrives. It was concluded that both methods can still give good results for estimating the significant wave height but not only Barrick’s method remains relevant for the estimation of the peak wave period in the omnidirectional case. The first trials of the FOBP method were presented in this paper with promising results. The technique will be pursued and hopefully improved in future works.
@@ Acknowledgment
Part of this research and the first author (VMM) have been funded by the Agence Nationale de la Recherche (ANR) under grant ANR-22-ASTR-0006-01 (ROSMED project: “Radar à Ondes de Surface en MEDiterranée”). We acknowledge the MOOSE program (Mediterranean Ocean Observing System for the Environment) coordinated by CNRS-INSU and the Research Infrastructure ILICO (CNRS-IFREMER). We thank the Parc National de Port-Cros (PNPC), the “Association Syndicale des Propriétaires du Cap Bénat” (ASPCB) as well as the Group Military Conservation (GMC) and the Marine Nationale for hosting our radar installation in Porquerolles, Cap Bénat and Fort Peyras, respectively. In addition, we would like to thank Alejandro Cáceres-Euse for providing the WWIII simulation results in the study region, which were used as a complementary dataset to validate the methodology here presented.

### v2

@@ B M W B M W B M W
Figure 11: Scatter plot of radar $H_{s}$ and $f_{p}$ from Modulation method versus CANDHIS buoy where each point is colored by the spatial density of nearby points.
4.2 Global comparison with WWIII
Although the WWIII model is not a perfect ground truth, it can be used for a coarse evaluation of the radar estimates over the entire radar coverage. For each radar cell and the entire time series, we calculated the RMS difference between the $H_{s}$ predicted by WWIII and the $H_{s}$ estimated by either Barrick’s or the Modulation method as well as the Pearson correlation coefficient between these quantities. The time coverage was also calculated as the proportion of successful estimates in the time series (after elimination of the outliers). Note that the tests used to select the second-order peaks in Barrick’s method results in a lower time coverage for this latter method.
Figure 12 shows the maps obtained for the statistical indices using the two methods, with the calibration factor $0.58$ applied to Barrick’s method estimate of $H_{s}$. The values considered to be irrelevant have been eliminated and shown in white. These are the radar cells for which the time coverage is lower than $30\%$ or the Pearson correlation coefficient is lower than $0.3$. The map confirms that Barrick’s method is more accurate than the Modulation method in estimating the $H_{s}$, with a RMSD of the order of $40$ cm in a large central zone (and probably better if the inaccuracy of WWIII itself is accounted for) out to about $30$ km range. The Pearson correlation coefficient remains larger than about $0.8$ in the same area. The Modulation method has in general a larger RMSD (of the order of $60$ cm at best) and smaller Pearson correlation coefficient (less than $0.7$). However, it provides coarse estimates with a high temporal coverage on a large distance. The same conclusions apply to the peak wave frequency (Figure 13) but with smaller values of the Pearson correlation coefficient for the two methods. Note that the modulation method is strongly affected around some ranges (number 15, 25-30 corresponding to distances of about 22 and 37-45 km) by the power grid artifact mentioned in Figure 6. However, relevant estimates are obtained in the far central zone and will now be discussed in more details.
Figure 12: Spatial maps of the statistical performance indicators for the radar measurement of $H_{s}$ obtained for the year $2020$ with the calibrated Barrick’s method (left column) and the modulation method (second column) when the WWIII is taken as reference: RMSD (in meter)(top); Pearson correlation coefficient (middle); Time coverage (in percent) (bottom). The polar grid has steps of 5 units in range (7.5 km) and 5 degrees in azimuth. The black arcs mark the ranges r5, r20, and r35 ($8$, $30$, and $52.5$ km) respectively; the black radial indicates the azimuth $78$ of the CANDHIS buoy used for the calculations in Table 2 while the magenta indicates azimuth $40$ used for the long range calculation (Section 4.3 below).
Figure 13: Same as Figure 12 for the peak wave frequency (in Hz).
4.3 Estimation in the far range
As mentioned above, it is interesting to examine the robustness of the two methods with respect to the range. To do this, we calculated the statistical performance indices along two distinct radials. The first one is the azimuth $78$ of the CANDHIS buoy, which is marked with a thick black line in Figure 12; the second is azimuth $40$, which is directed towards the open sea and is marked with a thick magenta line. Figure 14 shows the evolution of the RMSD, Pearson correlation coefficient and time coverage along these two azimuths for the Barrick’s and Modulation method, where WWIII is again taken as the reference. As can be seen, the Modulation method remains relevant at long range in the open sea, with a Pearson correlation coefficient greater than $0.5$ even though the RMSD becomes large. This is visually confirmed by the time series of $H_{s}$ and $f_{p}$ shown in Figure 15 at the cell range $30$ ($45$ km). While Barrick’s method yields only sparse estimates, the Modulation method is able to follow the trends of the wave parameters, albeit with sometimes large errors. In addition, Barrick’s method fails in estimating the small or moderate values of SWH (less than about $1$ m) while the Modulation method remains sensitive to the variations of the SWH even for small values. In order to understand this difference, we compared the Doppler spectrum and the Modulation spectrum obtained over on hour at this range cell (Figure 16). While the second-order Doppler spectrum is no longer distinguishable, the amplitude modulation of the analytical signal is still visible and the corresponding power spectrum has a clear peak at a dominant frequency. This explains why this last method can provide relevant (if not accurate) estimates of the wave parameters out to a farther distance than Barrick’s method.
Figure 14: Statistical performance indicators for Barrick’s method (blue lines) and the Modulation method (red lines). The black lines represent the RMS values of $H_{s}$ and $f_{p}$ according to the WWIII model at each range cell. The thick lines correspond to the azimuth $78$ (CANDHIS) while the thin line correspond to the azimuth $40$ (open sea).
Figure 15: Time series of $H_{s}$ (top) and $f_{p}$ (bottom) in the long range ($45$ km) along the azimuth $40$ (magenta line of Figure 12) as estimated by the WWIII model (red lines), Barrick’s method (blue symbols) or the Modulation method (magenta symbols).
Figure 16: a) Example of a Doppler spectrum obtained over one hour in the long range ($45km$); b) Visualization of the amplitude modulation through the absolute analytical signal; c) Power spectrum of the absolute analytical signal.
4.4 Discussion
Barrick’s method, which is based on the second-order Doppler spectrum, has well-known limitations, which we summarize below. The roughness parameter $K_{0}H_{s}$, where $K_{0}$ is the electromagnetic wave number, should not be too large for otherwise the perturbation expansions in the scattering orders would break down. This is clearly satisfied in the present case since at $16.15$ MHz with the local wave conditions ($H_{s}<4$ m), we have $K_{0}H_{s}<1.2$ Another condition for the validity of the asymptotic formulae 5 and 7 (see Barrick, 1977a ) is that the peak wave frequency of the ocean wave spectrum be much lower than the Bragg frequency, that is $f_{p}<<f_{B}$, a condition that can be violated at low sea states. This explains why the quality of the estimation deteriorates as the mean frequency increases, as has already been mentioned in Section 4.1. In addition, the signal-to-noise ratio of the second-order Doppler spectrum increases with the sea state and may be insufficient in the case of a calm sea. For these reasons, Barrick’s method is not adapted for measuring small values of $H_{s}$ nor large values of $f_{p}$, a situation which in general corresponds to low wind speed. Beyond these conditions, Barrick’s approach is obviously limited in range because it requires the Doppler spectrum side lobes to be larger than the noise level. As a result, the estimation of wave parameters also has a shorter range than the estimation of current, which relies on the sole first-order Bragg peak. The limitations of the Modulation method are less obvious to establish. However, by construction this method also requires a clear separation of scales between the Bragg wave and the dominant ocean waves, which again excludes low sea states and peak wave frequencies near the Bragg frequency ($2.5$ seconds). The heuristic justification of the long-wave modulation is limited to small long wave steepness but does not a priori impose any condition on the roughness parameter $K_{0}H_{s}$. It is therefore expected to be robust at large wave heights. As it uses the full received voltage and not only its second-order part, the method is also expected to be able to cover a greater range than the Doppler-based method, as long as the slow amplitude modulation is visible. Since the latter is more pronounced when longer waves are more developed and have larger height and periods, the modulation method is expected to be favoured by high winds and to work better in the open sea than in coastal aeras. This is partly confirmed by our experimental study. However, some experimental artifacts (which are specific to our installation) such as the power grid instability, prevent us from using the full potential of the method and the full range of distances.
@@ 4.5 Comparison with the literature
A number of studies have been published in the literature on the estimation of peak wave heights and frequency from a single HFR station and a single radar frequency. Some of them reported systematic comparisons with in situ instruments over a significantly long period (at least several months) using either phased array (Wyatt et al.,, 2003; Ramos et al.,, 2009; Gomez et al.,, 2015; Lopez et al.,, 2016; Lopez and Conley,, 2019; Alattabi et al.,, 2019; Al-Attabi et al.,, 2021) or compact radar systems (Roarty et al.,, 2015; Zhou and Wen,, 2015; Long et al.,, 2011; Atan et al.,, 2016; Saviano et al.,, 2020; Basañez et al.,, 2020). All of these methods are based on the exploitation of the second-order Doppler spectrum using either Barrick’s original methods, the full wave spectrum inversion method, or empirical methods developed by the radar manufacturers. The performance obtained varies considerably depending on the geographical location, the radar frequency, the measurement range, the antenna array, the inversion method chosen, the data post-processing choices, etc. However, two general trends can be identified within this diversity of results. First, the estimation of $H_{s}$ with phased array systems is generally slightly better than that with compact systems, with high Pearson correlation coefficients ($>0.9$) and RMSD often in the order of 30-40 cm and less than about 60 cm in all studies. The observed RMSD with compact systems are generally larger, in the order of 40-80 cm, with a large scatter of values and in some cases very large RMSD ($>1$ m); the associated Pearson correlation coefficients are generally smaller, mostly in the range of 0.5-0.9, except for the study Atan et al., (2016), which reports exceptionally good performances of 18-29 cm RMSD and 0.78-0.86 Pearson correlation coefficient. However, the decision to eliminate the $>25$% outliers (when compared with buoy measurement) before calculating the statistics is questionable. The second point is that the estimation of the peak frequency is less systematic and generally of lower quality than that of $H_{s}$, with Pearson correlation coefficients of the order of 0.5-0.7.
Our results are in agreement with these trends. The RMSD and Pearson correlation coefficient obtained for the estimation of $H_{s}$ with Barrick ranges from $0.3$ to about $0.5$ m and $0.8$ to $0.9$, respectively, for a wide angular sector in the near and medium range (up to about 35 km, see Figure 12c). The estimation of $f_{p}$ is much coarser, with a lower Pearson correlation coefficient ($0.6-0.7$ and an RMSD of at best $0.045$ (that is a relative error of the order of $20$%). However, these performances are based on a comparison with the WWIII which implies an extra error. The actual performances when compared with actual in situ measurements would be probably better.
@@ 5 Conclusion
In this work, we have tested the accuracy of the classical Barrick’s method for estimating the main sea state parameters based on weighted integrals of the second-order backscattered Doppler spectrum. For this purpose, we analyzed a one-year HFR data set from a monostatic station in the Toulon region. The results presented in this study have been compared with the data collected by the CANDHIS buoy located near the island of Porquerolles as well as hourly data simulated from the high-resolution WWIII model. For the classical Barrick’s method, the results obtained for the SWH have a good accuracy compared to the time series of the CANDHIS buoy, with a very high correlation and a low RMSD value between both of them (close to $90\%$ and around $0.3$ m, respectively, at $30$ km from the radar). A scaling factor of $0.58$ was found to be necessary in order to adjust the statistics of the two data sets and to further reduce the RMSD. With regard to the estimation of the peak wave frequency, this method has some shortcomings in capturing high values, due to insufficient scale separation between the long wave and the Bragg wave.
In addition, a new method, called the “Bragg Wave Modulation Method” (the “Modulation Method” for short), was also introduced and evaluated together with Barrick’s method. This technique uses the amplitude modulation of the received voltage time series and does not assume a decomposition of the signal in perturbation orders. It was found to be much less accurate than Barrick’s method at this stage but has been found to be promising for estimating wave characteristics at greater distances from the radars. A further limitation of this new method, at least for the $H_{s}$, is that it requires a preliminary calibration procedure with some ground truth measurements. This is not the case with Barrick’s method, which is based on the ratio of the first and second-order Doppler spectra and is therefore insensitive to absolute calibration (although a rather universal scaling factor improves its accuracy). The “Modulation Method” will be pursued and hopefully improved in future works.
@@ Acknowledgment
Part of this research and the first author (VMM) have been funded by the Agence Nationale de la Recherche (ANR) under grant ANR-22-ASTR-0006-01 (ROSMED project: “Radar à Ondes de Surface en MEDiterranée”). We acknowledge the MOOSE program (Mediterranean Ocean Observing System for the Environment) coordinated by CNRS-INSU and the Research Infrastructure ILICO (CNRS-IFREMER). We thank the Parc National de Port-Cros (PNPC), the “Association Syndicale des Propriétaires du Cap Bénat” (ASPCB) as well as the Group Military Conservation (GMC) and the Marine Nationale for hosting our radar installation in Porquerolles, Cap Bénat and Fort Peyras, respectively. In addition, we would like to thank Alejandro Cáceres-Euse for providing the WWIII simulation results in the study region.